16. Use incremental analysis to select the best alternatives using Benefit- Cost ration 4000 2000 60001000 9000 10000 7330 Cost Pw(benefits) 47000 1340 9500 Useful life is 20years Interest 6%
Answer #1

Incremental Analysis D& B

D

B

incremental cost/benefit

cost

1000

2000

1000

benefits

2340

47000

44660

incremental cost benefit ratio = incremental benefit/incremental cost = 44660/1000

44.66

Project D would be eliminated

Incremental analysis B & A

B

A

incremental cost/benefit

cost

2000

4000

2000

benefits

47000

7330

-39670

incremental cost benefit ratio = incremental benefit/incremental cost = -39670/2000

-19.835

Project A is eliminated

Incremental analysis B & C

B

A

incremental cost/benefit

cost

2000

6000

4000

benefits

47000

8730

-38270

incremental cost benefit ratio = incremental benefit/incremental cost = -38270/4000

-9.5675

Project C would be eliminated

Incremental analysis B & E

B

A

incremental cost/benefit

cost

2000

9000

7000

benefits

47000

9000

-38000

incremental cost benefit ratio = incremental benefit/incremental cost = -38000/7000

-5.4285714

Project E would be eliminated

Incremental analysis B & F

B

A

incremental cost/benefit

cost

2000

10000

8000

benefits

47000

9500

-37500

incremental cost benefit ratio = incremental benefit/incremental cost = -37500/8000

-4.6875

Project F would be eliminated

So on the basis of cost benefit analysis Project B would be selected